# Beyond Pixels: Enhancing LIME with Hierarchical Features and Segmentation Foundation Models

### v4

@@ C.2 Explaining wrong classification
Figure 10 shows an image of a hybrid between a horse (sorrel) and a zebra, where EfficientNetB4 can recognize both animals but does not contain the hybrid class. We explore why EfficientNetB4 assigns the highest probability to the zebra class rather than the sorrel. Although this is not strictly a misclassification, it simulates a similar situation and provides insight into why the model favors the zebra label over the sorrel. This analysis helps us understand the model’s decision-making process in cases where it prioritizes specific features associated with one class over another.
C.3 Stability of explanations
@@
The stability of imagery explanations using LIME can be linked to the quality of feature segments, as illustrated in Figure 11. This figure presents the segmentation maps generated by various techniques alongside their explanations and coefficient distributions, displayed through an IQR plot over eight runs. Notably, the DSEG technique divides the image into meaningful segments; the gorilla segment, as predicted by the EfficientNetB4 model, is distinctly visible and sharply defined. In contrast, other techniques also identify the gorilla, but less clearly, showing significant variance in their coefficient distributions. Watershed, while more stable than others, achieves this through overly broad segmentation, creating many large and a few small segments. These findings align with our quantitative evaluation and the described experimental setup.
C.4 Zero-shot classification explanation
In this section, we demonstrate the versatility of DSEG-LIME by applying it to a different dataset and classification task. Specifically, we replicate the zero-shot classification approach described in (Prasse et al., 2023) using CLIP (Radford et al., 2021) for the animal super-category. Since DSEG-LIME maintains model-agnostic properties, it remains applicable to zero- and few-shot classification models without modification.
Figure 12 presents an illustrative example from the dataset, where the task is to classify an image into the animal category. The predicted and ground-truth class for the image is ’Land mammal’. As shown by DSEG-LIME’s explanation, the model’s decision is primarily influenced by the presence of a deer in the foreground and a mountain in the background, which contribute to the overall classification.
Figure 12: DSEG-LIME explanation for CLIP. This figure illustrates an image processed by CLIP for zero-shot classification into animal categories. The model correctly predicted the class as ”Land mammal.” DSEG-LIME highlights the two most important features influencing the classification, with the presence of the deer being the most significant.
C.5 Exemplary limitation of DSEG
The example in Figure 13 shows a complex case of a hermit crab in front of sand, which is hardly detectable. Here, SAM fails to segment the image into meaningful segments, a known issue in the community (Khani et al., 2024). In contrast, SLIC can generate segments; thus, LIME can produce an explanation that does not show a complete image.
@@ C.5 Exemplary limitation of DSEG
C.6 Feature attribution maps
In addition to visualizing the $n$ most essential segments for an explanation, feature attribution maps also help the explainee (the person receiving the explanation (Miller, 2019)) to get an idea of which other segments are important for interpreting the result. In these maps, the segments represent the corresponding coefficient of the surrogate model learned within LIME for the specific case. Figure 14 represents all feature maps of EfficientNetB4 with the reported settings, accompanied by the original image with the most probable class. Blue segments are positively associated with the class to be explained, and red segments are negatively associated.
@@ D.1 Dataset
Image selection. As mentioned in Section 5.1, we selected various classes of images from the ImageNet (Deng et al., 2009) and COCO (Lin et al., 2014) dataset. Additionally, we created artificial images using the text-to-image model DALL-E (Ramesh et al., 2021) to challenge the XAI techniques when facing multiple objects. The dataset for the evaluation comprised 97 real images and three synthetic images. For the synthetic instances, the prompts ’realistic airplane at the airport’ (Airplane), ’realistic person running in the park’ (Miniskirt), and ’realistic person in the kitchen in front of a dishwasher’ (Dishwasher) were used.
The object types listed in Table 19 represent the primary labels of the images used in the dataset. Each image is unique, ensuring no duplication and maximizing the diversity of animals and objects covered. The bolded types denote those that were randomly selected for qualitative evaluation. The cursive types signify the reduced dataset comprising fifty instances, implemented to conserve computational time for supplementary evaluations. These types provide a balanced representation of the dataset and were chosen to ensure broad coverage across different categories. This selection strategy helps to avoid bias and supports a comprehensive evaluation.
@@ D.2 User study
Explanations. In Figures 17 and 17 we show all 20 images from the dataset used for the qualitative evaluation. Each image is accompanied by the prediction of EfficientNetB4 and the explanations within the vanilla LIME framework with all four segmentation approaches and the DSEG variant. The segments shown in the image indicate the positive features of the explanation.
@@
Instruction. Participants were tasked with the following question for each instance: ’Please arrange the provided images that best explain the concept [model’s prediction], ranking them from 1 (least effective) to 5 (most effective).’ Each instance was accompanied by DSEG, SLIC, Watershed, Quickshift, Felzenszwalb, and Watershed within the vanilla LIME framework and the hyperparameters discussed in the experimental setup. These are also the resulting explanations used in the quantitative evaluation of EfficientNetB4. Figure 18 shows an exemplary question of an instance of the user study.
Results. We show the cumulative maximum ratings in Figure 21 and in Figure 21 the median (in black), the interquartile range (1.5), and the mean (in red) for each segmentation technique. DSEG stands out in the absolute ratings, significantly exceeding the others. Similarly, in Figure 21, DSEG achieves the highest rating, indicating its superior performance relative to other explanations. Therefore, while DSEG is most frequently rated as the best, it consistently ranks high even when it is not the leading explanation, as the IQR of DSEG shows. Aligned with the quantitative results in Section 5.2, the Quickshift algorithm performs the worst.
@@
Table 20 presents the statistical significance of the user study. Specifically, it lists the t-statistics and p-values for comparisons between DSEG (the baseline method) and other segmentation methods, namely SLIC, QS, FS, and WS. The t-statistics indicate the magnitude of difference between DSEG and each different method, with higher values representing greater differences. The corresponding p-values demonstrate the probability that these observed differences are due to random chance, with lower values indicating stronger statistical significance.
@@
p-values $\downarrow$ 8.0e-143 $<$ 2.2e-308 1.2e-86 3.3e-187
In this context, the null hypothesis (H0) posits no significant difference in participant preferences between the performance of DSEG and other segmentation methods within the LIME framework. The alternative hypothesis (HA) asserts that DSEG performs significantly better than the different segmentation methods on the dataset when evaluated using five explanations. Given the extremely low p-values (e.g., 8.0e-143 for SLIC and ¡ 2.2e-308 for QS), we can reject the null hypothesis (H0) with high confidence. The significance level of 99.9% ($\alpha$ = 0.001) further supports this conclusion, as all p-values fall well below this threshold. These results indicate that the observed differences are highly unlikely to have occurred by chance and are statistically significant.

### v5

@@ C.3 Explaining wrong classification
Figure˜11 shows an image of a hybrid between a horse (sorrel) and a zebra, where EfficientNetB4 can recognize both animals but does not contain the hybrid class. We explore why EfficientNetB4 assigns the highest probability to the zebra class rather than the sorrel. Although this is not strictly a misclassification, it simulates a similar situation and provides insight into why the model favors the zebra label over the sorrel. This analysis helps us understand the model’s decision-making process in cases where it prioritizes specific features associated with one class over another.
C.4 Stability of explanations
@@
The stability of imagery explanations using LIME can be linked to the quality of feature segments, as illustrated in Figure˜12. This figure presents the segmentation maps generated by various techniques alongside their explanations and coefficient distributions, displayed through an IQR plot over eight runs. Notably, the DSEG technique divides the image into meaningful segments; the gorilla segment, as predicted by the EfficientNetB4 model, is distinctly visible and sharply defined. In contrast, other techniques also identify the gorilla, but less clearly, showing significant variance in their coefficient distributions. Watershed, while more stable than others, achieves this through overly broad segmentation, creating many large and a few small segments. These findings align with our quantitative evaluation and the described experimental setup.
C.5 Zero-shot classification explanation
In this section, we demonstrate the versatility of DSEG-LIME by applying it to a different dataset and classification task. Specifically, we replicate the zero-shot classification approach described in [43] using CLIP [44] for the animal super-category. Since DSEG-LIME maintains model-agnostic properties, it remains applicable to zero- and few-shot classification models without modification.
Figure˜13 presents an illustrative example from the dataset, where the task is to classify an image into the animal category. The predicted and ground-truth class for the image is ’Land mammal’. As shown by DSEG-LIME’s explanation, the model’s decision is primarily influenced by the presence of a deer in the foreground and a mountain in the background, which contribute to the overall classification.
Figure 13: DSEG-LIME explanation for CLIP. This figure illustrates an image processed by CLIP for zero-shot classification into animal categories. The model correctly predicted the class as "Land mammal." DSEG-LIME highlights the two most important features influencing the classification, with the presence of the deer being the most significant.
C.6 Exemplary limitation of DSEG
The example in Figure˜14 shows a complex case of a hermit crab in front of sand, which is hardly detectable. Here, SAM fails to segment the image into meaningful segments, a known issue in the community [26]. In contrast, SLIC can generate segments; thus, LIME can produce an explanation that does not show a complete image. However, to mitigate this issue, one could use a segmentation model that has better performance in such scenes, such as [10].
@@ C.6 Exemplary limitation of DSEG
C.7 Feature attribution maps
In addition to visualizing the $n$ most essential segments for an explanation, feature attribution maps also help the explainee (the person receiving the explanation [37]) to get an idea of which other segments are important for interpreting the result. In these maps, the segments represent the corresponding coefficient of the surrogate model learned within LIME for the specific case. Figure˜15 represents all feature maps of EfficientNetB4 with the reported settings, accompanied by the original image with the most probable class. Blue segments are positively associated with the class to be explained, and red segments are negatively associated.
@@ D.1 Dataset
Image selection. As mentioned in Section˜5.1, we selected various classes of images from the ImageNet [13] and COCO [31] dataset. Additionally, we created artificial images using the text-to-image model DALL-E [46] to challenge the XAI techniques when facing multiple objects. The dataset for the evaluation comprised 97 real images and three synthetic images. For the synthetic instances, the prompts ’realistic airplane at the airport’ (Airplane), ’realistic person running in the park’ (Miniskirt), and ’realistic person in the kitchen in front of a dishwasher’ (Dishwasher) were used.
The object types listed in Table˜19 represent the primary labels of the images used in the dataset. Each image is unique, ensuring no duplication and maximizing the diversity of animals and objects covered. The bolded types denote those that were randomly selected for qualitative evaluation. The cursive types signify the reduced dataset comprising fifty instances, implemented to conserve computational time for supplementary evaluations. These types provide a balanced representation of the dataset and were chosen to ensure broad coverage across different categories. This selection strategy helps to avoid bias and supports a comprehensive evaluation.
@@ D.2 User study
Explanations. In Figures˜18 and 18 we show all 20 images from the dataset used for the qualitative evaluation. Each image is accompanied by the prediction of EfficientNetB4 and the explanations within the vanilla LIME framework with all four segmentation approaches and the DSEG variant. The segments shown in the image indicate the positive features of the explanation.
@@
Instruction. Participants were tasked with the following question for each instance: ’Please arrange the provided images that best explain the concept [model’s prediction], ranking them from 1 (least effective) to 5 (most effective).’ Each instance was accompanied by DSEG, SLIC, Watershed, Quickshift, Felzenszwalb, and Watershed within the vanilla LIME framework and the hyperparameters discussed in the experimental setup. These are also the resulting explanations used in the quantitative evaluation of EfficientNetB4. Figure˜19 shows an exemplary question of an instance of the user study.
Results. We show the cumulative maximum ratings in Figure˜22 and in Figure˜22 the median (in black), the interquartile range (1.5), and the mean (in red) for each segmentation technique. DSEG stands out in the absolute ratings, significantly exceeding the others. Similarly, in Figure˜22, DSEG achieves the highest rating, indicating its superior performance relative to other explanations. Therefore, while DSEG is most frequently rated as the best, it consistently ranks high even when it is not the leading explanation, as the IQR of DSEG shows. Aligned with the quantitative results in Section˜5.2, the Quickshift algorithm performs the worst.
@@
Table˜20 presents the statistical significance of the user study. Specifically, it lists the t-statistics and p-values for comparisons between DSEG (the baseline method) and other segmentation methods, namely SLIC, QS, FS, and WS. The t-statistics indicate the magnitude of difference between DSEG and each different method, with higher values representing greater differences. The corresponding p-values demonstrate the probability that these observed differences are due to random chance, with lower values indicating stronger statistical significance.
@@
p-values $\downarrow$ 8.0e-143 << 2.2e-308 1.2e-86 3.3e-187
In this context, the null hypothesis (H0) posits no significant difference in participant preferences between the performance of DSEG and other segmentation methods within the LIME framework. The alternative hypothesis (HA) asserts that DSEG performs significantly better than the different segmentation methods on the dataset when evaluated using five explanations. Given the extremely low p-values (e.g., 8.0e-143 for SLIC and < 2.2e-308 for QS), we can reject the null hypothesis (H0) with high confidence. The significance level of 99.9% ($\alpha$ = 0.001) further supports this conclusion, as all p-values fall well below this threshold. These results indicate that the observed differences are highly unlikely to have occurred by chance and are statistically significant.